# Semiclassical Canovaccio for Composite Operators

### v1

@@ Appendix B Elliptic functions properties
$H(u\,|\,m)=\vartheta_{1}\left(\frac{\pi\,u}{2\,\mathbb{K}},q\right)\,,\qquad\Theta(u\,|\,m)=\vartheta_{4}\left(\frac{\pi\,u}{2\,\mathbb{K}},q\right)\,,\qquad Z(u\,|\,m)=\frac{\pi}{2\,\mathbb{K}}\,\frac{\vartheta_{4}^{\prime}(\frac{\pi\,u}{2\,\mathbb{K}},q)}{\vartheta_{4}(\frac{\pi\,u}{2\,\mathbb{K}},q)}\,$, (179)
@@
Lastly, using $Z(u|m)=E(u|m)-\mathbb{E}/\mathbb{K}\ F(u|m)$, one can obtain the useful integral representation
@@ C.1 $\phi^{4}$ theory
$2\pi(\sqrt{A_{\ell}}-1)-\frac{3\pi(A_{\ell}-2)}{2\sqrt{A_{\ell}}}m-\frac{3\pi\left(13A_{\ell}^{3}-49A_{\ell}^{2}+72A_{\ell}-24\right)}{32(A_{\ell}-1)A^{3/2}}m^{2}+\mathcal{O}\left({m^{3}}\right)\,.$ (193)

### v2

@@ Appendix B Elliptic functions properties
$H(u\,|\,m)=\vartheta_{1}\left(\frac{\pi\,u}{2\,{\mathbb{K}}(m)},q\right)\,,\qquad\Theta(u\,|\,m)=\vartheta_{4}\left(\frac{\pi\,u}{2\,{\mathbb{K}}(m)},q\right)\,,\qquad Z(u\,|\,m)=\frac{\pi}{2\,{\mathbb{K}}(m)}\,\frac{\vartheta_{4}^{\prime}(\frac{\pi\,u}{2\,{\mathbb{K}}(m)},q)}{\vartheta_{4}(\frac{\pi\,u}{2\,{\mathbb{K}}(m)},q)}\,$, (180)
@@
Lastly, using $Z(u|m)=E(u|m)-{\mathbb{E}}(m)/{\mathbb{K}}(m)\ F(u|m)$, one can obtain the useful integral representation
@@ C.1 $\phi^{4}$ theory
$2\pi(\sqrt{A_{\ell}}-1)-\frac{3\pi(A_{\ell}-2)}{2\sqrt{A_{\ell}}}m-\frac{3\pi\left(13A_{\ell}^{3}-49A_{\ell}^{2}+72A_{\ell}-24\right)}{32(A_{\ell}-1)A_{\ell}^{3/2}}m^{2}+\mathcal{O}\left({m^{3}}\right)\,.$ (194)